Respectfull tutor!!plz answer immediatly iam in hurry.Thankyou.The question is:James travel from Rawalpindi to lahore at a specific speed.He returned from there at a speed 5knm/h more than the speed he came.The distance between Rawalpindi to lahore is 380km.Total time taken by him to go and return is 8 hours.Find his speed both for going and returning.
Answer by scott8148(6628)

time = distance / rate

(380 / r) + (380 / (r + 5)) = 8

dividing by 4 ___ (95 / r) + (95 / (r + 5)) = 2

multiplying by r(r+5) ___ 95r + 475 + 95r = 2r^2 + 10r

subtracting 190r+475 ___ 0 = 2r^2 - 180r - 475

use quadratic formula to find r